A company purchased property for $100,000. The property included a building, a parking lot, and land. The building was appraised at $54,500; the land at $51,000, and the parking lot at $19,500. Land should be recorded in the accounting records with an allocated cost of:
Answer #1
Assets Fair value Percentage of fair value Allocated cost
Building 54,500 54,500/125,000 = 43.6% 100,000 x 43.6% = $43,600
Parking lot 19,500 19,500/125,000 = 15.6% 100,000 x 15.6% = $15,600
Land 51,000 51,000/125,000 = 40.8% 100,000 x 40.8% = $40,800
Total $125,000 100% $100,000

Land should be recorded at $40,800

kindly give a positive rating if you are satisfied with the solution. Please ask if you have any query related to the question, Thanks.